OpenStudy (anonymous):

What's the vertex of the parabola y=(x-5)^2+3

OpenStudy (anonymous):

-b/2a (x-5)^2+3 X^2-10x+25+3 x^2-10x+28 -(-10)/2(1)=5 = x y=(5-5)^2+3 y=3 Vertex at (5,3)